OpenStudy (anonymous):

A ladder 10 ft long rests against a vertical wall. If the bottom of the ladder slides away from the wall at a rate of 1.4 ft/s, how fast is the angle between the ladder and the ground changing when the bottom of the ladder is 6 ft from the wall?

OpenStudy (sleepyhead314):

calculus... related rates... I can probably help, gimme a sec to think :P

OpenStudy (sleepyhead314):



OpenStudy (anonymous):

L=10?

OpenStudy (sleepyhead314):

yes :) and dL/dt = 0 because the length of the ladder does not change over time

OpenStudy (sleepyhead314):

then they gave us dG/dt = 1.4 when g = 6 and they want us to find d(angle)/dt = ?

OpenStudy (sleepyhead314):

I think we can use cos for this

OpenStudy (anonymous):

I'm a little confused on how to set up the equation. Related Rates is my enemy.

OpenStudy (sleepyhead314):

it was mine too a year ago xD I just suck at trig right now... but I think it should go like cos(angle) = G/L take the derivative of that, do you know how?

OpenStudy (anonymous):

hmm, i think i can try fom here, give me a sec.

OpenStudy (sleepyhead314):

mhmm ^_^ chain rule on the left quociant on the right then once you get that, solve for d(angle)/dt

OpenStudy (anonymous):

nope, I'm stuck. I used the quotient for the right and got 14/10^2? and derivative of cos=-sin

OpenStudy (sleepyhead314):

don't plug in yet ^_^ so I got -sin(angle) * (d(angle)/dt) = ( L(dG/dt) - G(dL/dt) ) / L^2

OpenStudy (anonymous):

oh ok, so can i plug in the numbers now?

OpenStudy (anonymous):

i wish this were easier

OpenStudy (sleepyhead314):

lol me too mmm first get d(angle)/dt alone on one side and know that dL/dt = 0

OpenStudy (sleepyhead314):

sooo... d(angle)/dt = (dG/dt)/L * (-1/sin(angle)) now in order to plug, first find sin(angle) = ?

OpenStudy (sleepyhead314):

cow patties, I gtg sin(angle) = 8/10 dG/dt = 1.4 L = 10 plug it all in I got d(angle)/dt = -0.175 radians

OpenStudy (anonymous):

oh wow thank you!